OpenStudy (anonymous):

a candy store makes an 11- pound mixture of gummy candy, jelly beans, and hard candy. The cost of gummy candy is $2.00 per pound, jelly beans cost $3.00 per pound, and hard candy costs $3.00 per pound. The mixture calls for two times as many gummy candy pieces as jelly beans. The total cost of this mixture is $27.00. How much of each ingredient did the store use? 1. 6 lbs. gummy, 3 lbs. jelly beans, 2 lbs hard candy 2. 3 lbs. gummy candy, 2 lbs. jelly beans, 6 lbs hard candy 3. 6 lbs gummy, 2 lbs jelly beans, 3 lbs. hard candy 4. 2 lbs gummy candy, 3 lbs. jelly beans, 6 lbs hard candy

hero (hero):

g = pounds of gummy j = pounds of jelly h = pounds of hard g + j + h = 11 2g + 3j + 3h = 27

OpenStudy (anonymous):

Thank you. I did the math but I got it wrong tho. THANK YOU

hero (hero):

g = 2j

hero (hero):

g + j + h = 11 2g + 3j + 3h = 27 g = 2j 2j + j + h = 11 2(2j) + 3j + 3h = 27 3j + h = 11 7j + 3h = 27 9j + 3h = 33 7j + 3h = 27 Subtract the equations to get 2j = 6 j = 3 g = 6 h = 2

hero (hero):

Which means you should have picked the first answer choice.

OpenStudy (anonymous):

That the one I had. Ha. Thanks for your hard work.

OpenStudy (anonymous):

is it d? @Hero @Fez1370

OpenStudy (anonymous):

sorry. i dont remember

OpenStudy (anonymous):

its ok:) @Fez1370

OpenStudy (just_one_last_goodbye):

o_o @elliecosme ?